State true or false giving proper justification for following statement:there exists a continuous function f:RtoR2 such that f(down)=(n,1/n) for all n belongs to N

1
Expert's answer
2021-01-04T17:01:45-0500

"f:\R\rightarrow \R^2\\
f(cos(n))=(n,\frac{1}{n})"

Let f is continuous on "[-1,1]"

"\Rightarrow" "f" is bounded on "[-1,1]".

i.e.

"\exists M\in \N \ni \\
\;\;\;\;\;\;\; \;|f(x)-f(1)|<M\forall x\in [-1,1]"


Now take "x=cos (M+1)"

"\therefore f(x)=(M+1,\frac{1}{M+1})\\
\Rightarrow |f(x)-f(1)|=|(M+1,\frac{1}{M+1})-(1,1)|\\\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\sqrt{M^2+(1-(\frac{1}{M+1})^2}\\
\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;>\sqrt{M^2}=M\\\;\\
\therefore |f(x)-f(1)|>M"

That contradicts the bounded property.


Therefore "\pmb{\exists no\; f\ni f}" is continuous.